OpenStudy (anonymous):

what is the least common multiple of 450x^2y^6 and 3000x^4y^3

OpenStudy (zzr0ck3r):

hint: LCM(450,3000)=9000, LCM(x^2,x^4)=x^4 and LCM(y^6, y^3)=y^6